Diffusion Morning Report 04/11/11

Ted Berg tberg at stanford.edu
Mon Apr 11 09:37:10 PDT 2011


Good Morning All,
     Looking good this morning. Epi is down for HCl chaNGE. LTO has 
intermittent O2 flow issues. 4160 PM in progress.No TLC flow on Thermco 
2 No other issues at this time.

					
	Date: 	4/11/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*.88/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*.12/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*11.03/30* 	** 	**
Tylan 11-LTO 	*yellow* 	*O2 flow issues* 	*9.32/30* 	** 	*2*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*yellow* 	*loader issues* 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*down* 	*Hcl change* 	** 	** 	*1*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 469 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110411/e1fd7841/attachment.html>


More information about the diffusion mailing list